OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, -1). A. y = x - 3 B. y = x + 17 C. y = -3x - 3 D. y = -3x + 17

OpenStudy (campbell_st):

the line in point slope form is y = -3x + 7 The gradient(slope) of the line is -3 A perpendicular has a gradient of -1/m Using the data the slope of the perpendicular is -1/(-3) = 1/3 in point slops form the line is y = 1/3 x + b Use the point (6, - 1) to find b -1 = 1/3 (6) + b b = -3 the line is y = 1/3 x - 3 perhaps there is a typo in you choices...